# Another look at inference after prediction

### v1

@@ 5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\bm{\beta}^{*}_{1}$. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
@@
Figure 2 displays the CI length ratio across the different settings. In Scenario 1, we observe that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI lengths much less 1, indicating that the PB inference methods have reduced variance relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $\pi=n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, particularly when the labeled dataset size is larger, due to the difference in the augmentation terms of these two estimators. In this scenario, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
@@
Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscores our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average CI length ratio of 0.9 across the various sample sizes, indicating a gain in efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and have confidence intervals up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. This behavior is exacerbated in Scenario 2b, where the CI of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is up to 50% longer than the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and up to 72% longer for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$0.181 0.124 0.650 0.430 0.165
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$1.070 0.181 0.426 -0.423 -1.205
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$46.980 -47.523 -45.946 -46.725 -47.695
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$64.272 -64.115 -63.307 -63.759 -64.627
@@ 6 Discussion
In this paper, we showed that adding a simple weight to the PPI estimator provides valid inference and is guaranteed to be at least as efficient as estimation using only labeled data. The weight is drawn from the work of Chen and Chen (2000) in the context of double sampling. In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that can limit its practical utility. By incorporating the weight derived from Chen and Chen (2000), the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference. In addition to theoretical results that support these findings, our simulation studies demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate. Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide some interesting context for the PB inference problem. While we focused on least squares estimation, these results extend to other settings. The work of Chen and Chen (2000) and Angelopoulos et al. (2023a) apply to a wide variety of models beyond linear regression, such as parameters defining generalized linear models or parameters derived from estimating equations.
The work of Angelopoulos et al. (2023a) was pivotal in advancing the growing body of literature in PB inference. Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data Zrnic and Candès (2024b, b). Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches. The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator (Angelopoulos et al., 2023b). The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based solely on labeled data. In contrast, the approach of Chen and Chen (2000) minimizes the mean square error among a class of estimators with a linear augmentation term. Simultaneously, Miao et al. (2023) introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data. Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach. The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the augmentation term.
@@ A.1 Preliminaries
Notation. We denote the outer product of a vector $\boldsymbol{v}$ as $\boldsymbol{v}^{\otimes 2}=\boldsymbol{v}\boldsymbol{v}^{{\sf\scriptscriptstyle{T}}}$. We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix. We define the difference in asymptotic variance-covariance matrices between two estimators $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$ as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
@@ A.2 Proof of Theorem 1
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
@@
This inequality is strict provided that $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$. If the regression model is correctly specified (i.e., $\mathbb{E}(Y\mid{\bf{X}})={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$), then $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are asymptotically equivalent.
@@
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}.$
@@
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}},{\bf{X}})]=\bm{\beta}^{*}.$
@@ C.1 Proof of Theorem 2
Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.

### v2

@@ 5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\beta^{*}_{1}$. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
@@
Figure 2 displays the CI length ratio across the different settings. In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI length ratios below 1, indicating that the PB inference methods have reduced standard error relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $\pi=n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, particularly when the labeled dataset size is larger, due to the difference in the augmentation terms of these two estimators. In this scenario, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
@@
Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average CI length ratio of 0.9 across the various sample sizes, indicating a gain in efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes as there is not much to gain from the unlabeled data due to its relatively small size and the fact that $\widehat{Y}$ is a weak prediction of $Y$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios. In Scenario 2a, these estimators have CIs up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. This behavior is exacerbated in Scenario 2b, where the CI of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is up to 50% longer than the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and up to 72% longer for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$0.181 0.124 0.650 0.430 0.165
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$1.070 0.181 0.426 -0.423 -1.205
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$46.980 -47.523 -45.946 -46.725 -47.695
@@
     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$64.272 -64.115 -63.307 -63.759 -64.627
@@ 6 Discussion
In this paper, we showed that adding a simple weight to the PPI estimator provides valid PB inference that is guaranteed to be at least as efficient as estimation using only labeled data. The weight is drawn from the work of Chen and Chen (2000) in the context of double sampling. In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that can limit its practical utility. By incorporating the proposed weight, the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference. In addition to theoretical results that support these findings, our simulation studies demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate. Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide interesting context for the PB inference problem. While we focused on least squares estimation, our results extend to other settings. The work of Chen and Chen (2000) and Angelopoulos et al. (2023a) apply to a wide variety of settings beyond linear regression, such as generalized linear models.
The work of Angelopoulos et al. (2023a) was pivotal in advancing the growing body of literature in PB inference. Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data (Zrnic and Candès, 2024b, b). Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches. The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator (Angelopoulos et al., 2023b). The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based on labeled data. In contrast, the approach of Chen and Chen (2000) minimizes the mean square error among a class of estimators with a linear augmentation term. Simultaneously, Miao et al. (2023) introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data. Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach. The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the linear augmentation term.
@@ A.1 Preliminaries
Notation. We denote the outer product of a vector $\boldsymbol{v}$ as $\boldsymbol{v}^{\otimes 2}=\boldsymbol{v}\boldsymbol{v}^{{\sf\scriptscriptstyle{T}}}$. We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix. We define the difference in asymptotic variance-covariance matrices between two estimators, $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$, as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
@@ A.2 Proof of Theorem 1
$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
@@
This inequality is strict provided that $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$. If the regression model is also correctly specified (i.e., $\mathbb{E}(Y\mid{\bf{X}})={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$), then $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are asymptotically equivalent.
@@
$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}.$
@@
$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}})]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}},{\bf{X}})]=\bm{\beta}^{*}.$
@@ C.1 Proof of Theorem 2
Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.